# Cost-Effective Online Multi-LLM Selection with Versatile Reward Models

### v1

@@ A.1 Discussions on Constraint Type
A.2 Exploration on Attributes of Reward Functions
@@
❸ For the AIC reward, consider the vector $\boldsymbol{Z}\vcentcolon=(z_{1},\dots,z_{K})\in\mathbb{R}^{K}$, where $\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})=\exp\left(\sum_{k\in\mathcal{K}}z_{k}\ln\mu_{k}\right)$ is demonstrated to be a convex function with respect to $\boldsymbol{Z}$. This is evidenced by the Hessian matrix of $\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})$, denoted as $H_{\tilde{r}}(\boldsymbol{Z})\in\mathbb{R}^{K}\times\mathbb{R}^{K}$. The entries of $H_{\tilde{r}}(\boldsymbol{Z})$ are given by $H_{\tilde{r}}(\boldsymbol{Z})_{i,j}=\ln\mu_{i}\ln\mu_{j}\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})$. The Hessian can be expressed as $H_{\tilde{r}}(\boldsymbol{Z})=\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})\boldsymbol{W}\boldsymbol{W}^{T}$, where $\boldsymbol{W}$ is the column vector $\boldsymbol{W}\vcentcolon=(\ln\mu_{1},\dots,\ln\mu_{K})$. Consequently, $H_{\tilde{r}}(\boldsymbol{Z})$ is positive semi-definite because for any vector $\boldsymbol{x}\in\mathbb{R}^{K}$, we have $\boldsymbol{x}^{T}H_{\tilde{r}}(\boldsymbol{Z})\boldsymbol{x}=\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})(\boldsymbol{W}^{T}x)^{2}\geq 0$, given that $\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})\geq 0$ and $(\boldsymbol{W}^{T}\boldsymbol{x})^{2}\geq 0$. Applying Jensen’s inequality and acknowledging that $\mathbb{E}_{\boldsymbol{Z}}[z_{k}]=\tilde{z}_{k}$ as per Eq. 14, it follows that $\mathbb{E}_{\boldsymbol{Z}}[\exp\left(\sum_{kin\mathcal{K}}z_{k}\ln\mu_{k}\right)]\geq\exp\left(\sum_{kin\mathcal{K}}\mathbb{E}_{\boldsymbol{Z}}[z_{k}]\ln\mu_{k}\right)=\tilde{r}(\tilde{\boldsymbol{Z}},\boldsymbol{\mu})$, thereby ensuring that $\mathbb{E}\left[\tilde{r}\left(\mathbb{I}{S_{t}},\boldsymbol{\mu}\right)\right]\geq\tilde{r}(\tilde{\boldsymbol{Z}}_{t},\boldsymbol{\mu})$.
Appendix B Proof Appendix
B.1 Proof of Lemma 1
@@
B.2 Comprehensive Proof of Theorem 1
@@
$\mathbb{E}\left[r\left(S_{t},\bar{\boldsymbol{\mu}}_{t}\right)-r\left(S_{t},\boldsymbol{\mu}\right)\right]\leq\frac{1}{o^{*}}\mathbb{E}\left[\sum_{t=1}^{T}r\left(S_{t},\bar{\boldsymbol{\mu}}_{t}\right)-r\left(S_{t},\boldsymbol{\mu}\right)\mathbb{I}\left\{F_{t}=\left|S_{t}\right|\right\}\right]$ (17)
$\leq\frac{L}{o^{*}}\mathbb{E}\left[\sum_{t=1}^{T}\sum_{\begin{subarray}{c}k\in S_{t}\\ k\leq\mathcal{F}_{t}\end{subarray}}\left|\bar{\mu}_{t,k}-\mu_{t,k}\right|\right]\leq\frac{L}{o^{*}}\mathbb{E}\left[\sum_{t=1}^{T}\sum_{\begin{subarray}{c}k\in S_{t}\\ k\leq\mathcal{F}_{t}\end{subarray}}2\rho_{t,\mu_{k}}\right].$
@@
Define the over-estimation regret for C2MAB-V at round $t$ as $\bar{reg}(S_{t},\bar{\boldsymbol{\mu}}_{t})=r(S_{t},\bar{\boldsymbol{\mu}}_{t})-r(S_{t},\boldsymbol{\mu})$, where $S_{t}$ denotes the action and $\bar{\boldsymbol{\mu}}_{t}$ represents the upper confidence bound (UCB) value of base arms at round $t$. The history of C2MAB-V prior to selecting action $S_{t}$, denoted as $\mathcal{H}_{t}$, encompasses the sequence of actions and observations up to round $t-1$, expressed as $\mathcal{H}_{t}=((S_{1},\mathcal{F}_{1},\boldsymbol{\mu}_{1,\mathcal{F}_{1}},\boldsymbol{c}_{1,S_{1}}),\dots,$ $(S_{t-1},\mathcal{F}_{t-1},\boldsymbol{\mu}_{t-1,\mathcal{F}_{t-1}},c_{t-1,S_{t-1}}))$. Here, $\boldsymbol{\mu}_{t,\mathcal{F}_{t}}=(\boldsymbol{\mu}_{t,S_{t,1}},...,\boldsymbol{\mu}_{t,S_{t,\mathcal{F}_{t}}})$ and $\boldsymbol{c}_{t,S_{t}}=(c_{t,S_{t,1}},...,\boldsymbol{c}_{t,S_{t,|S_{t}|}})$ detail the partial observed rewards and full costs of LLM for the $t^{th}$ action, respectively.
@@
B.3 Comprehensive Proof of Theorem 2
@@
To address violation (a), we introduce a series of random variables $\{Y_{1},Y_{2},\dots,Y_{T}\}$, constructing a martingale sequence relative to the historical data of C2MAB-V, denoted as $\mathcal{H}_{t}$. The history $\mathcal{H}_{t}$ encompasses the sequence of actions and observed outcomes up to round $t-1$, encapsulating both the actions taken and the corresponding observed rewards and full costs. Formally, $\mathcal{H}_{t}$ includes tuples $((S_{1},\mathcal{F}_{1},\boldsymbol{\mu}_{1,\mathcal{F}_{1}},\boldsymbol{c}_{1,S_{1}}),\dots,$ $(S_{t-1},\mathcal{F}_{t-1},\boldsymbol{\mu}_{t-1,\mathcal{F}_{t-1}},c_{t-1,S_{t-1}}))$, where $\boldsymbol{\mu}_{t,\mathcal{F}_{t}}=(\boldsymbol{\mu}_{t,S_{t,1}},...,\boldsymbol{\mu}_{t,S_{t,\mathcal{F}_{t}}})$ and $\boldsymbol{c}_{t,S_{t}}=(c_{t,S_{t,1}},...,\boldsymbol{c}_{t,S_{t,|S_{t}|}})$, representing the observed rewards and costs, respectively, for the action set $S_{t}$.
@@
Similar to how we prove violation (a), we will define a series of random variables $\{Y_{1},\dots,Y_{T}\}$ to form a martingale with respect to history of C2MAB-V $\mathcal{H}_{t}$, defined as $\mathcal{H}_{t}=((S_{1},\mathcal{F}_{1},\boldsymbol{\mu}_{1,\mathcal{F}_{1}},\boldsymbol{c}_{1,S_{1}}),\dots,$ $(S_{t-1},\mathcal{F}_{t-1},\boldsymbol{\mu}_{t-1,\mathcal{F}_{t-1}},c_{t-1,S_{t-1}}))$ to encompasses the sequence of actions and observations up to round $t-1$, before choosing the action $S_{t}$ which includes the first $t-1$ actions and first $t-1$ partial observed rewards and full costs, where $\boldsymbol{\mu}_{t,\mathcal{F}_{t}}=(\boldsymbol{\mu}_{t,S_{t,1}},...,\boldsymbol{\mu}_{t,S_{t,\mathcal{F}_{t}}})$ and $\boldsymbol{c}_{t,S_{t}}=(c_{t,S_{t,1}},...,\boldsymbol{c}_{t,S_{t,|S_{t}|}})$. The martingale sequence is defined recursively starting with $\boldsymbol{Y}_{0}=0$. For each round $t=1,\dots,T$, the incremental change is given by:
@@ B.4 Technical Inequalities
Appendix C More Experiments
C.1 Respective Results for Reward and Violation
@@
C.2 In-depth Analysis of Exploratory Results.

### v2

@@ C.1 Discussions on Constraint Type and Extended Task Type
Based on the above analysis, regarding extended tasks, such as identifying the set of LLMs to be deployed, are addressed, the strategy lacks specificity on which LLM should be assigned to a particular task (as in the SUC case) or which LLM should handle a given sub-module (as in the AIC case). This issue can be mitigated by transitioning from cardinality-constrained subsets to partition matroids. Specifically, a smaller and more suitable subset of LLMs can be preselected for specific query tasks. For example, in mathematical problems, the selection space for LLMs could be restricted to those fine-tuned for mathematical queries, rather than choosing from the full set of LLMs. This targeted selection not only reduces the search space but also accelerates the online learning process by avoiding the misallocation of resources.
Alternatively, the selection process can rely entirely on online learning through match bandit. The match bandit problem, a variant of the maximum weight matching problem, is modeled as a bipartite graph $G=(Q,K,E)$, where $Q$ represents query tasks, $K$ represents LLMs, and $E$ represents weighted edges denoting the probability of assigning a query task $q$ to an LLM $k$. The Combinatorial MAB proposed in our paper can naturally be transformed into this bipartite graph structure based on .[45, 14].
C.2 Exploration on Attributes of Reward Functions
@@
❸ For the AIC reward, consider the vector $\boldsymbol{Z}\vcentcolon=(z_{1},\dots,z_{K})\in\mathbb{R}^{K}$, where $\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})=\exp\left(\sum_{k\in\mathcal{K}}z_{k}\ln\mu_{k}\right)$ is demonstrated to be a convex function with respect to $\boldsymbol{Z}$. This is evidenced by the Hessian matrix of $\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})$, denoted as $H_{\tilde{r}}(\boldsymbol{Z})\in\mathbb{R}^{K}\times\mathbb{R}^{K}$. The entries of $H_{\tilde{r}}(\boldsymbol{Z})$ are given by $H_{\tilde{r}}(\boldsymbol{Z})_{i,j}=\ln\mu_{i}\ln\mu_{j}\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})$. The Hessian can be expressed as $H_{\tilde{r}}(\boldsymbol{Z})=\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})\boldsymbol{W}\boldsymbol{W}^{T}$, where $\boldsymbol{W}$ is the column vector $\boldsymbol{W}\vcentcolon=(\ln\mu_{1},\dots,\ln\mu_{K})$. Consequently, $H_{\tilde{r}}(\boldsymbol{Z})$ is positive semi-definite because for any vector $\boldsymbol{x}\in\mathbb{R}^{K}$, we have $\boldsymbol{x}^{T}H_{\tilde{r}}(\boldsymbol{Z})\boldsymbol{x}=\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})(\boldsymbol{W}^{T}x)^{2}\geq 0$, given that $\tilde{r}(\boldsymbol{Z},\boldsymbol{\mu})\geq 0$ and $(\boldsymbol{W}^{T}\boldsymbol{x})^{2}\geq 0$. Applying Jensen’s inequality and acknowledging that $\mathbb{E}_{\boldsymbol{Z}}[z_{k}]=\tilde{z}_{k}$ as per Eq. 14, it follows that $\mathbb{E}_{\boldsymbol{Z}}[\exp\left(\sum_{k\in\mathcal{K}}z_{k}\ln\mu_{k}\right)]\geq\exp\left(\sum_{k\in\mathcal{K}}\mathbb{E}_{\boldsymbol{Z}}[z_{k}]\ln\mu_{k}\right)=\tilde{r}(\tilde{\boldsymbol{Z}},\boldsymbol{\mu})$, thereby ensuring that $\mathbb{E}\left[\tilde{r}\left(\mathbb{I}{S_{t}},\boldsymbol{\mu}\right)\right]\geq\tilde{r}(\tilde{\boldsymbol{Z}}_{t},\boldsymbol{\mu})$.
Appendix D Proof Appendix
D.1 Proof of Lemma 1
@@
D.2 Comprehensive Proof of Theorem 1
@@
$\quad\mathbb{E}\left[r\left(S_{t},\bar{\boldsymbol{\mu}}_{t}\right)-r\left(S_{t},\boldsymbol{\mu}\right)\right]$
$\leq\frac{1}{o^{*}}\mathbb{E}\left[\sum_{t=1}^{T}r\left(S_{t},\bar{\boldsymbol{\mu}}_{t}\right)-r\left(S_{t},\boldsymbol{\mu}\right)\mathbb{I}\left\{F_{t}=\left|S_{t}\right|\right\}\right]$
$\leq\frac{L}{o^{*}}\mathbb{E}\left[\sum_{t=1}^{T}\sum_{\begin{subarray}{c}k\in\mathcal{F}_{t}\end{subarray}}\left|\bar{\mu}_{t,k}-\mu_{t,k}\right|\right]$
$\leq\frac{L}{o^{*}}\mathbb{E}\left[\sum_{t=1}^{T}\sum_{\begin{subarray}{c}k\in\mathcal{F}_{t}\end{subarray}}2\rho_{t,\mu_{k}}\right].$
@@
Define the over-estimation regret for C2MAB-V at round $t$ as $\bar{reg}(S_{t},\bar{\boldsymbol{\mu}}_{t})=r(S_{t},\bar{\boldsymbol{\mu}}_{t})-r(S_{t},\boldsymbol{\mu})$, where $S_{t}$ denotes the action and $\bar{\boldsymbol{\mu}}_{t}$ represents the upper confidence bound (UCB) value of base arms at round $t$. The history of C2MAB-V prior to selecting action $S_{t}$, denoted as $\mathcal{H}_{t}$, encompasses the sequence of actions and observations up to round $t-1$, expressed as $\mathcal{H}_{t}=((S_{1},\mathcal{F}_{1},\boldsymbol{\mu}_{1,\mathcal{F}_{1}},\boldsymbol{c}_{1,S_{1}}),\dots,$ $(S_{t-1},\mathcal{F}_{t-1},\boldsymbol{\mu}_{t-1,\mathcal{F}_{t-1}},\boldsymbol{c}_{t-1,S_{t-1}}))$. Here, $\boldsymbol{\mu}_{t,\mathcal{F}_{t}}=(\boldsymbol{\mu}_{t,S_{t,1}},...,\boldsymbol{\mu}_{t,S_{t,F_{t}}})$ and $\boldsymbol{c}_{t,S_{t}}=(c_{t,S_{t,1}},...,c_{t,S_{t,|S_{t}|}})$ detail the partial observed rewards and full costs of LLM for the $t^{th}$ action, respectively.
@@
D.3 Comprehensive Proof of Theorem 2
@@
To address violation (a), we introduce a series of random variables $\{Y_{1},Y_{2},\dots,Y_{T}\}$, constructing a martingale sequence relative to the historical data of C2MAB-V, denoted $\mathcal{H}_{t}$. The history $\mathcal{H}_{t}$ encompasses the sequence of actions and observed outcomes up to round $t-1$, encapsulating both the actions taken and the corresponding observed rewards and full costs. Formally, $\mathcal{H}_{t}$ includes tuples $((S_{1},\mathcal{F}_{1},\boldsymbol{\mu}_{1,\mathcal{F}_{1}},\boldsymbol{c}_{1,S_{1}}),\dots,$ $(S_{t-1},\mathcal{F}_{t-1},\boldsymbol{\mu}_{t-1,\mathcal{F}_{t-1}},\boldsymbol{c}_{t-1,S_{t-1}}))$, where $\boldsymbol{\mu}_{t,\mathcal{F}_{t}}=(\boldsymbol{\mu}_{t,S_{t,1}},...,\boldsymbol{\mu}_{t,S_{t,F_{t}}})$ and $\boldsymbol{c}_{t,S_{t}}=(c_{t,S_{t,1}},...,c_{t,S_{t,|S_{t}|}})$, representing the observed rewards and costs, respectively, for the action set $S_{t}$.
@@
Similar to how we prove violation (a), we will define a series of random variables $\{Y_{1},\dots,Y_{T}\}$ to form a martingale with respect to history of C2MAB-V $\mathcal{H}_{t}$, defined as $\mathcal{H}_{t}=((S_{1},\mathcal{F}_{1},\boldsymbol{\mu}_{1,\mathcal{F}_{1}},\boldsymbol{c}_{1,\mathcal{F}_{1}}),\dots,$ $(S_{t-1},\mathcal{F}_{t-1},\boldsymbol{\mu}_{t-1,\mathcal{F}_{t-1}},c_{t-1,\mathcal{F}_{t-1}}))$ to encompasses the sequence of actions and observations up to round $t-1$, before choosing the action $S_{t}$ which includes the first $t-1$ actions and first $t-1$ partial observed rewards and full costs, where $\boldsymbol{\mu}_{t,\mathcal{F}_{t}}=(\boldsymbol{\mu}_{t,S_{t,1}},...,\boldsymbol{\mu}_{t,S_{t,F_{t}}})$ and $\boldsymbol{c}_{t,S_{t}}=(c_{t,S_{t,1}},...,\boldsymbol{c}_{t,S_{t,F_{t}}})$. The martingale sequence is defined recursively starting with $\boldsymbol{Y}_{0}=0$. For each round $t=1,\dots,T$, the incremental change is given by:
@@ D.4 Technical Inequalities
Appendix E More Experiments
E.1 Experimental Details
Consistent with Section 6, we conduct all experiments on a device equipped with an Intel Core i5-13600KF CPU @ 3.50GHz and 32 GB of memory. Gurobi, a powerful mathematical optimization solver that supports various programming models, is utilized to solve relaxed constraint optimization problems, employing version 11.0.1 [27]. Fig. 5 depicts a sample interaction involving large language models (LLMs) across various scientific disciplines—Biology, Chemistry, Geography, and Physics—utilizing the SciQ dataset [69]. Table 3 lists the LLMs used in the experiments (refer to Section 6), detailing each model’s name with parameters, cost per 1,000 tokens in USD, and size in gigabytes. The models range from smaller configurations, such as ChatGLM2-6B-32K, to larger ones like Llama 2-70B. Considering the variability in generating answers when using LLMs, as well as potential non-responses, we specify reward allocations based on different LLM return scenarios in addition to comparisons with the SciQ dataset’s labels. If the question is answered accurately and adheres to the prescribed prompt format, the reward is set at 0.5. In the event of an incorrect response, no reward is granted. However, if factors such as network congestion or potential issues with the LLM result in an empty response, a reward of 0.1 is allocated. Should the response adhere to the format guidelines yet fail to meet the specified requirements, a reward of 0.3 is assigned. We simulate user feedback for each round using the public SciQ dataset to function as a binary feedback [61], and our algorithm learns the performance of LLMs based on this feedback. Our entire framework is built on the popular bandit framework [39], which continuously updates and adjusts decisions based on a steady stream of data (i.e., user feedback) to address the classic exploration-exploitation dilemma regarding LLM performance and cost.
LLM_ID Model Name (with Parameters) Cost (USD/1k tokens) Size (GB)
1 ChatGLM2-6B-32K [52] 0.005 12.5
2 ChatGPT-3.5 [52] 0.02 /
3 Claude 2 [15] 0.08 /
4 ERNIE 3.5-8K [5] 0.015 /
5 Llama 2-7B [51] 0.005 12.6
6 Llama 2-13B [51] 0.008 24.3
7 Llama 2-70B [51] 0.05 128.3
8 Mixtral-8x7B-Instruct [50] 0.05 93.37
9 ChatGPT-4 [52] 0.12 /
E.2 Respective Results for Reward and Violation
(a) AWC
(b) SUC
(c) AIC
Figure 6: Reward of three task types with nine different LLMs.
(a) AWC
(b) SUC
(c) AIC
Figure 7: Violation of three task types with nine different LLMs.
Per-round Reward and Violation. Fig. 6 and Fig. 7 display the experimental rewards and violations for three different task types. In Fig. 6, C2MAB-V consistently outperforms or matches the $\epsilon$-Greedy baseline in terms of rewards per round. The parameters on $\alpha_{\mu}$ and $\alpha_{c}$ used are (0.3, 0.05), (1, 0.05), (0.3, 0.01), and (1, 0.01), respectively denoted as (a), (b), (c), and (d). Adjusting the values of $\alpha_{\mu}$ and $\alpha_{c}$ consistently results in higher rewards for C2MAB-V settings. These settings also significantly reduce violations per round compared to the $\epsilon$-Greedy approach. Although the CUCB algorithm achieves the highest per-round rewards, it does so completely ignoring cost constraints, leading to a per-round violation rate at least four times higher than that of C2MAB-V, as shown in Fig. 7. This figure also indicates that C2MAB-V’s violations are lower than those of the $\epsilon$-Greedy baseline, highlighting its ability to effectively manage constraints. Additionally, since Thompson Sampling relies on Bayesian updates, which can be highly variable, this method may lack stability and potentially cause severe oscillations, as observed in scenarios like the AIC task.
Additionally, as shown in Fig. 6(c), we can observe a convergence trend in the rewards of C2MAB-V. This trend suggests that the initial actions selected might have had higher rewards but also higher violations. Over time, the cost-aware online learning strategy adjusts by slightly reducing rewards to achieve greater cost savings, demonstrating its effectiveness in managing long-term constraints.
@@ E.2 Respective Results for Reward and Violation
E.3 In-depth Analysis of Exploratory Results.
@@
Enhancing LLM Selection: From Two-Tier to Multi-Tier. We further demonstrate the benefits of transitioning from a traditional “two-tier” to a more complex “multi-tier” LLM selection strategy in online settings. In the simpler two-tier system, selections are limited to just two levels of LLM engagement, which might not adequately address the diverse range of queries typically encountered in dynamic environments. By adopting a multi-tier strategy, our system can more precisely match the complexity and specificity of incoming queries with an appropriate level of computational resources and LLM expertise. The ratio comparison is not displayed in Fig. 12, as the denominator may be zero. Fig. 12 underscores the necessity and effectiveness of our advanced multi-tier strategy in enhancing response quality and system adaptability in real-time online interactions.
(a) Reward
(b) Violation
Figure 12: Comparison with a “two-tier LLM selection strategy”, which only considers one larger and one smaller LLM, against a multi-tier approach under the AWC task type.
@@
Figure 13: Comparison with “offline-learned fixed sets of multi-LLM cascaded combinations” for “online queries” under the AWC task type.
Necessity of Online Learning for LLM Selection. As discussed in Section 6, our approach emphasizes online learning methodologies that incorporate continuous feedback. To highlight the significance of ongoing online adjustments in LLM selection, we present a comparison. Initially, we pre-learned a fixed combination of multiple LLMs offline, which was then utilized to manage online queries. As illustrated in Fig. 13, our feedback-driven adjustments significantly enhance the effectiveness of pre-set LLM selections. By comparing our method with purely offline-learned, cascaded combinations of multi-LLMs for the AWC task type, we demonstrate the essential role of continuous online adaptations in optimizing LLM selection for real-time queries.
(a) Reward
(b) Violation
(c) Ratio
Figure 14: Comparison on asynchronous local-cloud architecture.